OpenStudy (anonymous):

Short probability question: I may have gotten "lucky", but what are the chances that if I roll 2 (fair) 6-sided die, that I would correctly guess both numbers three out of four times?

OpenStudy (anonymous):

same idea to @UnkleRhaukus Qn last night :)

ganeshie8 (ganeshie8):

binomial distribution wid p = 1/36, n = 4, k = 3 : \[\large \binom{4}{3} \left(\frac{1}{36}\right )^3 \left(\frac{35}{36}\right)^1 \]

OpenStudy (anonymous):

Had to post this question after watching Through the Wormhole, is luck real :)

ganeshie8 (ganeshie8):

that number looks very unlikely ^

OpenStudy (anonymous):

luck is real ;)

OpenStudy (anonymous):

0.0000277841284

OpenStudy (anonymous):

very lucky to me lol

OpenStudy (anonymous):

or 35/1259712 thank you

OpenStudy (anonymous):

Another question, that is related

OpenStudy (anonymous):

what a nice display pic ! i sow something like this in a song but i forget what it is xD

ganeshie8 (ganeshie8):

check ur calculation once

OpenStudy (anonymous):

Oh I did (4/3)*(1/36)^3 * (35/6), not ( (4/3)*(1/36)^3 ) * (35/36)

ganeshie8 (ganeshie8):

\[\large \binom{4}{3} = ^4C_3 = 4\]

ganeshie8 (ganeshie8):

its a combination symbol :)

OpenStudy (anonymous):

Oh ok lol

OpenStudy (anonymous):

If you did it with C, I would of remembered

OpenStudy (anonymous):

Would the prob. change if I guessed the two numbers correctly 3 out of the 4 times, but the numbers for each die was switched? Like: I roll die 1 and die 2. I guess die 1 to be 1, and die 2 to be 2. The numbers are still 1 and 2, but die 2 is 1, and die 1 is 2

OpenStudy (anonymous):

*It would change, so:

OpenStudy (anonymous):

Would would be the prob. that I roll 2 fair 6-sided die, and guessed the numbers of each pair of die correctly 3 out of 4 times?

OpenStudy (anonymous):

The first question, either die could be one of the two numbers. What would be the prob. that each die I guessed would be specifically one of the two numbers?

OpenStudy (anonymous):

still here?

ganeshie8 (ganeshie8):

So now the order of prediction doesnt matter is it ?

ganeshie8 (ganeshie8):

if u predict (1, 2) then both below give u success : (1, 2) and (2, 1)

ganeshie8 (ganeshie8):

Is that the right ?

OpenStudy (anonymous):

Yes. So instead of saying what are the chances of (1,2) or (2,1); (3,4) or (4,3), and (5,6) or (6,5)... what are the chances of die (*1,2), (3,4), and then (5,6)

ganeshie8 (ganeshie8):

just find the probability for success (p) and plug that in the binomial formula ?

OpenStudy (anonymous):

ok, but I haven't taken statistics in years. This is just a random thing I wanted to figure out :)

ganeshie8 (ganeshie8):

we can work it using combinatorics

ganeshie8 (ganeshie8):

total equivalent outcomes = 6(6+1)/2 = 21 so, p = 1/21

ganeshie8 (ganeshie8):

the probability for predicting 3 out of 4 times correctly, when order of dies doesnt matter is : \[ \large \binom{4}{3} \left(\frac{1}{21}\right )^3 \left(\frac{20}{21}\right)^1 \]

ganeshie8 (ganeshie8):

let me knw if the number 21 is throwing u off :)

OpenStudy (anonymous):

so 80/194481

OpenStudy (anonymous):

I remember some things about stat, just not too much :)

OpenStudy (anonymous):

So the first answer you gave was when order mattered?

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

for the first problem : (1, 2) is different from (2, 1) so u get 36 total outcomes

ganeshie8 (ganeshie8):

So the chances of predicting correctly one time is 1/36

ganeshie8 (ganeshie8):

for ur second problem, the chance of predicting correctly one time is 1/21

OpenStudy (anonymous):

I understand a little more now. thanks again. ...Since when I did do it 2 years ago, when the order was messed up...I wasn't all that lucky.

ganeshie8 (ganeshie8):

haha cool :) btw that derivation of binomial formula is trivial - it directly follows from permutations

ganeshie8 (ganeshie8):

look it up in google when u r free...

OpenStudy (anonymous):

ok, thanks!

OpenStudy (anonymous):

Always good to get a refresher

ganeshie8 (ganeshie8):

this textbook is a real nice refresher on statistics : http://www.openintro.org/stat/textbook.php

ganeshie8 (ganeshie8):

click on "Download the Second Edition PDF" Enjoy...

ganeshie8 (ganeshie8):

`chapter 3.4` covers the binomial distribution